Ball A is thrown vertically up from the ground at time t=0 with an initial velocity of 60m/sec. Ball B is thrown vertically up from ground at t=0 with an initial velocity of 45m/sec. Calculate the maximum distance between the two balls and the time (t) at which this occurs.

Grade:12th pass

1 Answers

Eshan
5 years ago
From the frame of ball B, ball A moves with up with constant speed of 60m/s-45m/s=15m/s upwards with no relative acceleration. So the distance will constantly increase until the ball B reaches the ground.
Time of flight of ball B=\dfrac{2u_B}{g}=\dfrac{2\times 45}{10}=9s
Hence the maximum distance between the two balls in the time=v_{rel}\times t=15m/s\times 9=135m